# Boltzmanngasse 5, A 1090 Vienna, Austria

### v2

@@ Abstract
We construct, in even spacetime dimensions, a family of singularity-free Kerr–Anti-de Sitter-like black holes with negatively curved cross-sections of conformal infinity and non-spherical cross-sections of horizons.
@@ 1 Introduction
The Schwarzschild metrics, together with their higher-dimensional generalisations, are expected to exhaust the collection of $(n+1)$-dimensional, vacuum, static, asymptotically flat metrics satisfying natural global properties ChCo . [^2] See (ChodoshEtAl, , Remark 1.4 in v2) and references therein concerning the positive energy theorem; that last theorem is used in the proof of uniqueness. These metrics are a special case of the Birmingham-Kottler metrics Birmingham ; Kottler , which allow for any value of the cosmological constant $\Lambda$, and where the usual $(n-1)$-dimensional spherical sections of the event horizon can be replaced by any $(n-1)$-dimensional Einstein geometry.
@@
The Kerr metric is the unique $(3+1)$-dimensional vacuum stationary analytic and asymptotically flat black hole satisfying a set of reasonable properties, and its higher dimensional generalisation are likewise of interest. The family of higher dimensional, asymptotically flat Kerr-like vacuum metrics was introduced by Myers and Perry in Myers:1986un while the case in which we are interested, with non-zero cosmological constant, has been presented in Gibbons:2004uw ; Gibbons:2004js . In $(3+1)$ dimensions the algebraically special property of the Weyl tensor of the Kerr metric is often emphasised, and such algebraically special metrics have been considered and classified in higher dimensions, see Ortaggio:2017ydo ; Reall:2011ys and references therein; some partial uniqueness results have been obtained Lucietti:2020ltw ; Hollands:2012xy , see also MarsKdS for a uniqueness result when $\Lambda>0$. The Kerr-like class of metrics has been generalized to include multiple NUT parameters Chen:2006ea . We refer the interested reader to the review Emparan:2008eg for more examples of higher dimensional black holes.
@@
Now, the sections of the horizons of the metrics of Gibbons:2004uw ; Gibbons:2004js have spherical topology, which begs the question of existence of Kerr-like $(n+1)$-dimensional vacuum metrics with more general topologies of the horizon. The aim of this paper is to point out the existence of such metrics in the case of a negative cosmological constant, which are stationary outside of an event horizon, and have non-spherical sections of the horizon, in all even spacetime dimensions $(n+1)\geq 4$.
@@
The metrics we obtain are parameterised by a mass parameter $\bar{m}$ and by $N:=(n-1)/2$ rotation parameters $\bar{a}_{i}$. They generalise the metrics presented in Klemm:1997ea ; Klemm:1998kd , which have only one rotation parameter. They have the remarkable property of having no curvature singularities. Again remarkably, no causality violations arise when
@@
This work is organised as follows: In Section 2 we recall the higher-dimensional Kerr-type metrics with a cosmological constant. In Section 3 we present our new metrics. We show in Section 3.1 that the metrics reduce to the Birmingham-Kottler metrics when the rotation parameters vanish. In Sections 3.2-3.3 we establish extendibility across Killing horizons by rewriting the metric in a Kerr-Schild form, and show nonexistence of “non-coordinate” singularities. In Section 3.4 we construct a conformal completion at infinity. The signature of the metric is checked in Section 3.7. We give a sharp criterion on the parameters for absence of causality violations in Section 3.8, see Proposition 3.4 there. We show that the geometry of sections of the horizon cannot be realized on a compact manifold for small rotation parameters in Section 3.9; whether or not this can be done in general remains to be seen. Existence of black hole regions is established in Section 3.10. In Section 4 we construct projection diagrams for a class of natural maximal analytic extensions. The question of generalisation of the construction to other dimensions, positive cosmological constant, or Ricci flat sections of the horizon is briefly addressed in Section 5. Some technical results, in particular the analysis of the number of zeros of the equation determining the location of Killing horizons, are left to Appendices A-C.
@@ 2 Spherical Kerr-Anti-de Sitter black holes in higher dimensions
3 Multiply rotating Kerr–AdS-type black holes with non-spherical horizons
@@
which resembles how one would embed an $N$-dimensional hyperbolic space in one higher dimension. To retain the signature (,+,,+)(-,+$\ldots$,+)( - , + , … , + ) of the metric, we make the further substitutions
@@
Choose a mass parameter $\bar{m}\in\mathbb{R}$ and a cosmological constant parameter $\ell>0$. Consider any connected component, say ${\mycal U}$, of the set on which $\bar{g}$ is defined and non-degenerate. Proposition 3.1 below shows that such sets do not depend upon the rotation parameters, hence all the metrics $\bar{g}$ with the same value of $\bar{m}$ are simultaneously defined there. The calculations in Hamamoto:2006zf , together with the usual analytic continuation arguments, show that the tensor field $\bar{g}$ defines on ${\mycal U}$ a pseudo-Riemannian metric which solves the vacuum Einstein equations with cosmological constant
$\Lambda=-\frac{1}{2}n(n-1)\ell^{-2}\,,$ (3.6)
@@ 3.1 Non-rotating case
This is the Minkowski metric in $12N+$ 1 dimensions, which induces the hyperbolic metric on the spacelike hyperboloid (3.15).
@@
Happily, we have recovered the Birmingham–Kottler metric,
$g_{\mathrm{BK}}=f_{{{\varepsilon}}}d\bar{t}^{2}+\frac{d\bar{r}^{2}}{f_{{{\varepsilon}}}}+\bar{r}^{2}{\mathring{\gamma}}_{AB}dx^{A}dx^{B}\,,\quad f_{{{\varepsilon}}}={{\varepsilon}}+\ell^{-2}\bar{r}^{2}-2m\bar{r}^{n-2}\,.$ (3.18)
@@
In (3.18) the tensor field ${\mathring{\gamma}}_{AB}$ is an $(n-1)$-dimensional Einstein metric, and ${{\varepsilon}}$ is the sign of its scalar curvature. This also shows by continuity that the scalar curvature of the sections of the Killing horizons will be negative for small rotation parameters. However, as in the previously known cases Klemm:1997ea ; Klemm:1998kd it will not be constant.
@@ 3.2 Singularities
If $\bar{m}=0$ the metric is locally Minkowski/Anti-de Sitter in unusual coordinates Gibbons:2004uw , and we will demonstrate this explicitly in Section 3.3.
@@
The remaining potential singularities of the tensor field $\bar{g}$ arise from zeros of the determinant, of the $\bar{\mu}^{i}$’s, and of $\bar{V}-2\bar{m}\bar{r}$. It turns out that all these can be gotten rid of by introducing new coordinates.
Indeed, the zeros of the determinant are determined in Appendix A, with the results there summarised in Proposition 3.1 below.
The zeros of $\bar{\mu}^{i}$’s have already been shown to be innocuous polar-coordinates singularities at axes of rotation (cf. below (3.11)), and do not require further attention.
@@ 3.3 The Kerr–Schild extension
Similarly to the Myers–Perry metrics, the metric (3) can brought into Kerr–Schild form using a similar construction as in Gibbons:2004uw .
@@
One can use the $(\tau,\bar{r},\varphi^{i},\mu^{i})$-coordinates to express the metric (3) as
@@
$k_{\mu}dx^{\mu}$ $=\bar{W}d\tau+Fd\bar{r}+\left(\sum_{i=1}^{N}\frac{\bar{a}_{i}(\bar{\mu}^{i})^{2}}{\bar{\Xi}_{i}}d\varphi^{i}\right)\,,$ (3.30)
@@ 3.5 Extensions, Killing horizons
We wish to show that, in the fully-rotating case, the tensor field (3) can be extended through all the remaining potentially singular sets, namely zeros of $1-\ell^{-2}\bar{r}^{2}$ in the Kerr-Schild form of the metric, and zeros of $\bar{V}-2\bar{m}\bar{r}$ in (3).
@@ 3.5.1 $1-\ell^{-2}\bar{r}^{2}$
Possible concerns about the signature of (3) at zeros of $1-\ell^{-2}\bar{r}^{2}$ are dispelled by Proposition 3.1. So, as already mentioned, the original form of the metric provides an extension of the Kerr-Schild form of the metric across $\{\bar{r}=\pm\ell\}$.
@@ 3.5.2 $\bar{V}-2\bar{m}\bar{r}$
When $\bar{a}_{i}\neq 0$, the case $\bar{m}\bar{r}=0$ in (3.39) is only possible if $\bar{m}=0$. In this case the rotation parameters are actually irrelevant in that they can be gotten rid of by a coordinate transformation: Indeed, it follows from Section 3.3 that the metric $\bar{g}$ is then locally isometric to the Anti-de Sitter metric. Changing $\bar{r}$ to its negative if necessary, it is standard (cf., e.g., (ChBlackHoles, , Chapter 6)) to construct an extension where $\bar{r}>0$, and in which the hypersurface $\{\bar{r}=\ell\}$ is a bifurcate Killing horizon.
In the fully-rotating case (3.37), the roots of $\bar{V}-2\bar{m}\bar{r}$ occur away from the set $\{\bar{r}\pm\ell\}$. It then follows that the Kerr-Schild form of the metric (3.29)-(3.30) provides an analytic extension across each of the roots of $\bar{V}-2\bar{m}\bar{r}$. Further extensions are constructed in Section 4.
@@
Figure 3.1: Examples of the function $F$ with three (top) or four (bottom) zeros. In all plots, $\ell=1$ (black dashed line).
@@ 3.6 Surface gravity
±$\pm\kappa=\bar{r}_{H}(\ell^{-2}\bar{r}_{H}^{2}-1)\left(\sum_{i=1}^{N}\frac{1}{\bar{r}_{H}^{2}+\bar{a}_{i}^{2}}+\frac{1}{2\bar{r}_{H}^{2}}\right)+\frac{1}{\bar{r}_{H}}\,;$ (3.44)
@@ 3.7 Signature
Now, the analysis of Section 3.5.2 shows that the signature does not change across the zeros of $\bar{V}-2\bar{m}\bar{r}$, when $\bar{g}$ is extended across these zeros using the Kerr-Schild form of the metric. Further, it follows from (3.27) that $\bar{r}$ is a smooth function on the Kerr-Schild extension. Its level sets away from the zeros of $\bar{V}-2\bar{m}\bar{r}$ are all diffeomorphic to $\mathbb{R}^{2N+1}$ on the manifold underlying the original form (3) of the metric. Furthermore, (3.27) shows that
@@
Hence the level sets of $\bar{r}$ are also diffeomorphic to $\mathbb{R}^{2N+1}$ near the zeros of $\bar{V}-2\bar{m}\bar{r}$ in the Kerr-Schild extension.
We conclude that the extension of the metric (3) through each connected component of the set $\{\bar{V}-2\bar{m}\bar{r}=0\}$, as obtained by using the Kerr-Schild coordinates, provides a smooth Lorentzian metric defined on
@@
The function $\bar{r}$ is a globally defined smooth function on ${\mycal M}_{\mathrm{KS}}$, while the coordinates $(\bar{t},x^{i},y^{i})\in\mathbb{R}^{2N+1}$, where $(x^{i},y^{i})$ are as in (3.11), are smooth coordinates away from the zeros of $\bar{V}-2\bar{m}\bar{r}$, and have to be patched together using Kerr-Schild coordinates across the horizons.
@@ 3.8.1 Stable causality
which shows that each connected component of the set $\{\bar{V}-2\bar{m}\bar{r}<0\}$ is stably causal, with either $\bar{r}$ or $-\bar{r}$ being time functions there.
@@
$h:=h_{ij}d\phi^{i}d\phi^{j}=$ $\Big{(}\frac{\bar{r}^{2}+\bar{a}_{i}^{2}}{\bar{\Xi}_{i}}(\bar{\mu}^{i})^{2}\delta_{ij}+\frac{2\bar{m}\bar{r}\bar{a}_{i}\bar{a}_{j}(\bar{\mu}^{i})^{2}(\bar{\mu}^{j})^{2}}{\bar{U}\bar{\Xi}_{i}\bar{\Xi}_{j}}\Big{)}d\phi^{i}d\phi^{j}\,,$ (3.48)
$N^{j}=$ ${2\bar{m}\bar{r}}\bar{W}\left({\bar{U}}+{2\bar{m}\bar{r}}\,\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{(\bar{r}^{2}+\bar{a}_{i}^{2})\bar{\Xi}_{i}}\right)^{-1}\frac{\bar{a}_{j}}{\bar{r}^{2}+\bar{a}_{j}^{2}}\,,$ (3.49)
±$\pm N^{2}=$ $-(\bar{V}-2\bar{m}\bar{r})\frac{\bar{W}}{\left(1+\frac{2\bar{m}\bar{r}}{\bar{U}}\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{(\bar{r}^{2}+\bar{a}_{i}^{2})\bar{\Xi}_{i}}\right)\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})}\,,$ (3.50)
@@
$\bar{g}^{\bar{t}\,\bar{t}}$ =±$=\pm N^{-2}=-\frac{\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})}{\bar{W}(\bar{V}-2\bar{m}\bar{r})}\Bigg{(}1+\frac{2\bar{m}\bar{r}}{\bar{U}}\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{\Xi_{i}(\bar{r}^{2}+\bar{a}_{i}^{2})}\Bigg{)}$ (3.51)
@@
Since $\bar{W}>1$ and $\bar{U}>0$, Equation (3.51) shows that ¯$\bar{\nabla}\bar{t}$ is manifestly timelike on the set
@@
    At any point with $\bar{r}\neq 0$ we can make (3.54) negative while maintaining (3.8.1) by making $\bar{m}$ negative large enough. Thus there exist sets of parameters and subsets of $\{\bar{V}-2\bar{m}\bar{r}>0\}$ where ¯$\bar{\nabla}\bar{t}$ is not timelike. We will show in Section 3.8.2 that stable causality is violated whenever ()(*)( ∗ ) fails to be positive.

### v3

@@ Abstract
We construct, in even spacetime dimensions, a family of singularity-free Kerr-anti-de Sitter-like black holes with negatively curved cross-sections of conformal infinity and non-compact, non-spherical cross-sections of horizons, with the full set of rotation parameters.
@@ 1 Introduction
The Schwarzschild metrics, together with their higher-dimensional generalisations, are expected to exhaust the collection of $(n+1)$-dimensional, vacuum, static, asymptotically flat metrics satisfying natural global properties ChCo . [^3] See (ChodoshEtAl, , Remark 1.4 in v2) and references therein concerning the positive energy theorem; that last theorem is used in the proof of uniqueness. These metrics are a special case of the Birmingham-Kottler metrics Birmingham ; Kottler , which allow for any value of the cosmological constant $\Lambda$, and where the usual (n1)(n-1)-dimensional spherical sections of the event horizon can be replaced by any (n1)(n-1)-dimensional Einstein geometry.
@@
The Kerr metric is the unique $(3+1)$-dimensional vacuum stationary analytic and asymptotically flat black hole satisfying a set of reasonable properties, and its higher dimensional generalisation are likewise of interest. The family of higher dimensional, asymptotically flat Kerr-like vacuum metrics was introduced by Myers and Perry in Myers:1986un while the case in which we are interested, with non-zero cosmological constant, has been presented in Gibbons:2004uw ; Gibbons:2004js . In $(3+1)$ dimensions the algebraically special property of the Weyl tensor of the Kerr metric is often emphasised, and such algebraically special metrics have been considered and classified in higher dimensions, see Ortaggio:2017ydo ; Reall:2011ys and references therein; some partial uniqueness results have been obtained Lucietti:2020ltw ; Hollands:2012xy , see also MarsKdS for a uniqueness result when $\Lambda>0$. The Kerr-like class of metrics has been generalised to include multiple NUT parameters Chen:2006ea . We refer the interested reader to the review Emparan:2008eg for more examples of higher dimensional black holes.
@@
Now, the sections of the horizons of the metrics of Gibbons:2004uw ; Gibbons:2004js have spherical topology. A more general class of Kerr-AdS-like metrics, with negatively curved horizons and with a single rotation parameter has been discovered by Klemm, Moretti, and Vanzo Klemm:1997ea ; Klemm:1998kd . This begs question of existence of Kerr-like $(n+1)$-dimensional vacuum metrics with negatively curved horizons and with the whole set of rotation parameters. The aim of this paper is to point out the existence of such stationary metrics, in the case of a negative cosmological constant, with non-spherical sections of the horizon, in all even spacetime dimensions $(n+1)\geq 4$.
@@
The metrics we obtain are parameterised by a mass parameter $\bar{m}$ and by N:=(n1)/2N:=(n-1)/2 rotation parameters $\bar{a}_{i}$. As already pointed out, they generalise the metrics presented in Klemm:1997ea ; Klemm:1998kd , which have only one rotation parameter. In particular, when $\bar{m}\bar{a}_{1}...\bar{a}_{N}\neq 0$ the metrics have the remarkable property of having no curvature singularities. We have not investigated the question, whether or not the spacetimes are geodesically complete. It should be clear from the projection diagrams that we construct in Section 4 that the maximal extensions considered there are not globally hyperbolic; whether or not the spacetimes are geodesically complete, and whether the domains of outer communication are globally hyperbolic, would require further investigations. Again remarkably, no causality violations arise when
@@
This work is organised as follows: In Section 2 we recall the higher-dimensional Kerr-type metrics with a cosmological constant. In Section 3 we present our new metrics. We show in Section 3.1 that the metrics reduce to the Birmingham-Kottler metrics when the rotation parameters vanish. In Sections 3.2-3.3 we establish extendibility across Killing horizons by rewriting the metric in a Kerr–Schild form, and show nonexistence of “non-coordinate” singularities. In Section 3.4 we construct a conformal completion at infinity. The signature of the metric is checked in Section 3.7. We give a sharp criterion on the parameters for absence of causality violations in Section 3.8, see Proposition 3.4 there. We show that the geometry of sections of the horizon cannot be realised on a compact manifold for small rotation parameters in Section 3.9; whether or not this can be done in general remains to be seen. Existence of black hole regions is established in Section 3.10. In Section 4 we construct projection diagrams for a class of natural maximal analytic extensions. The question of generalisation of the construction to other dimensions, positive cosmological constant, or Ricci flat sections of the horizon is briefly addressed in Section 5. Some technical results, in particular the analysis of the number of zeros of the equation determining the location of Killing horizons, are left to Appendices A-C.
@@ 2 Spherical Kerr-Anti-de Sitter black holes in higher dimensions
3 Multiply rotating Kerr-AdS-type black holes with non-spherical horizons
@@
which resembles how one would embed an $N$-dimensional hyperbolic space in one higher dimension. To retain the signature (,+,,+)(-$\ldots,+$) of the metric, we make the further substitutions
@@
Choose a mass parameter $\bar{m}\in\mathbb{R}$ and a cosmological constant parameter $\ell>0$. Consider any connected component, say $\mycal U$}, of the set on which $\bar{g}$ is defined and non-degenerate. Proposition 3.1 below shows that such sets do not depend upon the rotation parameters, hence all the metrics $\bar{g}$ with the same value of $\bar{m}$ are simultaneously defined there. The calculations in Hamamoto:2006zf (which take place in a different coordinate system to those used here; a translation between the systems can be found in Sections 4.2-5 of Frolov:2017kze ) demonstrate that (2) satisfies the Einstein equations in all dimensions. This, together with the usual analytic continuation arguments (see Appendix D where they are spelled-out in detail), shows that the tensor field $\bar{g}$ defines on $\mycal U$} a pseudo-Riemannian metric which solves the vacuum Einstein equations with cosmological constant
$\Lambda=-\frac{1}{2}n(n-1)\ell^{-2}\,.$ (3.6)
@@ 3.1 Non-rotating case
This is the Minkowski metric in $2N+1$ dimensions, which induces the hyperbolic metric on the spacelike hyperboloid (3.15).
@@
Happily, we have recovered the Birmingham–Kottler metric Birmingham ; Kottler ,
$g_{\mathrm{BK}}=-f_{{{\varepsilon}}}d\bar{t}^{2}+\frac{d\bar{r}^{2}}{f_{{{\varepsilon}}}}+\bar{r}^{2}{\mathring{\gamma}}_{AB}dx^{A}dx^{B}\,,\quad f_{{{\varepsilon}}}={{\varepsilon}}+\ell^{-2}\bar{r}^{2}-2m\bar{r}^{n-2}\,.$ (3.18)
@@
In (3.18) the tensor field ${\mathring{\gamma}}_{AB}$ is an (n1)(n-1)-dimensional Einstein metric, and $\varepsilon$}} is the sign of its scalar curvature. This also shows by continuity that the scalar curvature of the sections of the Killing horizons will be negative for small rotation parameters. However, as in the previously known cases Klemm:1997ea ; Klemm:1998kd it will not be constant.
@@ 3.2 Singularities
If $\bar{m}=0$ the metric is locally Anti-de Sitter in unusual coordinates Gibbons:2004uw , and we will demonstrate this explicitly in Section 3.3.
@@
The remaining potential singularities of the tensor field $\bar{g}$ arise from zeros of the determinant, of the $\bar{\mu}^{i}$’s, and of $\bar{V}-2\bar{m}\bar{r}$. It turns out that all these can be got rid of by introducing new coordinates. Indeed, the zeros of the determinant are determined in Appendix A, with the results there summarised in Proposition 3.1 below. The zeros of $\bar{\mu}^{i}$’s have already been shown to be innocuous polar-coordinates singularities at axes of rotation (cf. below (3.11)), and do not require further attention.
@@ 3.3 The Kerr–Schild extension
Similarly to the Myers–Perry metrics, the metric (3) can be brought into Kerr–Schild form using a similar construction as in Gibbons:2004uw .
@@
One can use the $(\tau,\bar{r},\varphi^{i},\hat{\mu}^{i})$-coordinates to express the metric (3) as
@@
$k_{\mu}dx^{\mu}$ $=\bar{W}d\tau+Fd\bar{r}+\left(\sum_{i=1}^{N}\frac{\bar{a}_{i}(\hat{\mu}^{i})^{2}}{\bar{\Xi}_{i}}d\varphi^{i}\right)\,$, (3.30)
@@ 3.5 Extensions, Killing horizons
We wish to show that, in the fully-rotating case, the tensor field (3) can be extended through all the remaining potentially singular sets, namely zeros of $1-\ell^{-2}\bar{r}^{2}$ in the Kerr–Schild form of the metric, and zeros of $\bar{V}-2\bar{m}\bar{r}$ in (3).
@@ 3.5.1 $1-\ell^{-2}\bar{r}^{2}$
Possible concerns about the signature of (3) at zeros of $1-\ell^{-2}\bar{r}^{2}$ are dispelled by Proposition 3.1. So, as already mentioned, the original form of the metric provides an extension of the Kerr–Schild form of the metric across $\{\bar{r}=\pm\ell\}$.
@@ 3.5.2 $\bar{V}-2\bar{m}\bar{r}$
When $\bar{a}_{i}\neq 0$, the case $\bar{m}\bar{r}=0$ in (3.39) is only possible if $\bar{m}=0$. In this case the rotation parameters are actually irrelevant in that they can be got rid of by a coordinate transformation: Indeed, it follows from Section 3.3 that the metric $\bar{g}$ is then locally isometric to the Anti-de Sitter metric. Changing $\bar{r}$ to its negative if necessary, it is standard (cf., e.g., (ChBlackHoles, , Chapter 6)) to construct an extension where $\bar{r}>0$, and in which the hypersurface $\{\bar{r}=\ell\}$ is a bifurcate Killing horizon.
In the fully-rotating case (3.37), the roots of $\bar{V}-2\bar{m}\bar{r}$ occur away from the set $\{\bar{r}=\pm\ell\}$. It then follows that the Kerr–Schild form of the metric (3.29)-(3.30) provides an analytic extension across each of the roots of $\bar{V}-2\bar{m}\bar{r}$. Further extensions are constructed in Section 4.
@@
Figure 3.1: Examples of the function $F$ with three (top) or four (bottom) zeros. In all plots, $\ell=1$.
@@ 3.6 Surface gravity
$\pm\kappa=\bar{r}_{H}(\ell^{-2}\bar{r}_{H}^{2}-1)\left(\sum_{i=1}^{N}\frac{1}{\bar{r}_{H}^{2}+\bar{a}_{i}^{2}}+\frac{1}{2\bar{r}_{H}^{2}}\right)+\frac{1}{\bar{r}_{H}}\,;$ (3.44)
@@ 3.7 Signature
Now, the analysis of Section 3.5.2 shows that the signature does not change across the zeros of $\bar{V}-2\bar{m}\bar{r}$, when $\bar{g}$ is extended across these zeros using the Kerr–Schild form of the metric. Further, it follows from (3.27) that $\bar{r}$ is a smooth function on the Kerr–Schild extension. Its level sets away from the zeros of $\bar{V}-2\bar{m}\bar{r}$ are all diffeomorphic to $\mathbb{R}^{2N+1}$ on the manifold underlying the original form (3) of the metric. Furthermore, (3.27) shows that
@@
Hence the level sets of $\bar{r}$ are also diffeomorphic to $\mathbb{R}^{2N+1}$ near the zeros of $\bar{V}-2\bar{m}\bar{r}$ in the Kerr–Schild extension.
We conclude that the extension of the metric (3) through each connected component of the set $\{\bar{V}-2\bar{m}\bar{r}=0\}$, as obtained by using the Kerr–Schild coordinates, provides a smooth Lorentzian metric defined on
@@
The function $\bar{r}$ is a globally defined smooth function on ${\mycal M}_{\mathrm{KS}}$, while the coordinates $(\bar{t},x^{i},y^{i})\in\mathbb{R}^{2N+1}$, where $(x^{i},y^{i})$ are as in (3.11), are smooth coordinates away from the zeros of $\bar{V}-2\bar{m}\bar{r}$, and have to be patched together using Kerr–Schild coordinates across the horizons.
@@ 3.8.1 Stable causality
which shows that each connected component of the set $\{\bar{V}-2\bar{m}\bar{r}<0\}$ is stably causal, with either $\bar{r}$ or $\bar{r}$ being time functions there.
@@
$h:=h_{ij}d\phi^{i}d\phi^{j}=\$$\Big(\frac{\bar{r}^{2}+\bar{a}_{i}^{2}}{\bar{\Xi}_{i}}(\bar{\mu}^{i})^{2}\delta_{ij}+\frac{2\bar{m}\bar{r}\bar{a}_{i}\bar{a}_{j}(\bar{\mu}^{i})^{2}(\bar{\mu}^{j})^{2}}{\bar{U}\bar{\Xi}_{i}\bar{\Xi}_{j}}\Big)d\phi^{i}d\phi^{j}\,$, (3.48)
$N^{j}=\$${2\bar{m}\bar{r}}\bar{W}\left({\bar{U}}+{2\bar{m}\bar{r}}\,\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{(\bar{r}^{2}+\bar{a}_{i}^{2})\bar{\Xi}_{i}}\right)^{-1}\frac{\bar{a}_{j}}{\bar{r}^{2}+\bar{a}_{j}^{2}}\,$, (3.49)
$\pm N^{2}=\$$-(\bar{V}-2\bar{m}\bar{r})\frac{\bar{W}}{\left(1+\frac{2\bar{m}\bar{r}}{\bar{U}}\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{(\bar{r}^{2}+\bar{a}_{i}^{2})\bar{\Xi}_{i}}\right)\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})}\,$, (3.50)
@@
$\bar{g}^{\bar{t}\,\bar{t}}$ $=\pm N^{-2}=-\frac{\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})}{\bar{W}(\bar{V}-2\bar{m}\bar{r})}\Bigg(1+\frac{2\bar{m}\bar{r}}{\bar{U}}\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{\Xi_{i}(\bar{r}^{2}+\bar{a}_{i}^{2})}\Bigg)$ (3.51)
@@
Since $\bar{W}>1$ and $\bar{U}>0$, Equation (3.51) shows that $\bar{\nabla}\bar{t}$ is manifestly timelike on the set
@@
    At any point with $\bar{r}\neq 0$ we can make (3.54) negative while maintaining (3.8.1) by making $\bar{m}$ negative large enough. Thus there exist sets of parameters and subsets of $\{\bar{V}-2\bar{m}\bar{r}>0\}$ where $\bar{\nabla}\bar{t}$ is not timelike. We will show in Section 3.8.2 that stable causality is violated whenever ()(*) fails to be positive.